OpenStudy (nickymail48):

(x-2)(x+6)=9 how to solve

OpenStudy (blank ):

You need step by step instructions or answer?

OpenStudy (nickymail48):

instructions plz

OpenStudy (blank ):

Just a second...

OpenStudy (blank ):

(x-2)(x+6)=9 Multiply each term in the first group by each term in the second group using the FOIL method. FOIL stands for First Outer Inner Last, and is a method of multiplying two binomials. First, multiply the first two terms in each binomial group. Next, multiply the outer terms in each group, followed by the inner terms. Finally, multiply the last two terms in each group. (x*x+x*6-2*x-2*6)=9 Simplify the FOIL expression by multiplying and combining all like terms. x^(2)+4x-12=9 To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side. x^(2)+4x-21=0 For a polynomial of the form x^(2)+bx+c, find two factors of c (-21) that add up to b (4). In this problem 7*-3=-21 and 7-3=4, so insert 7 as the right hand term of one factor and -3 as the right-hand term of the other factor. (x+7)(x-3)=0 Set each of the factors of the left-hand side of the equation equal to 0. x+7=0 x-3=0 Since 7 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 7 from both sides. x=-7 x-3=0 Set each of the factors of the left-hand side of the equation equal to 0. x=-7 x-3=0 Since -3 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 3 to both sides. x=-7 x=3 The complete solution is the set of the individual solutions. x=-7,3 Hope this helps. :)

OpenStudy (nickymail48):

thanks

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)